# Dolphin v1.0 Technical Report

### v2

@@ 2.1.4 Distilled knowledge data
Following the methodology of (Xu et al., 2023b), we first constructed a set of ultrasound data question templates. The difficulty level of questions was enhanced through depth-wise augmentation, while their diversity was improved via breadth-wise expansion. All question categories were strictly standardized within the ultrasound medical domain, and we distilled ultrasound-specific data from both GPT and DeepSeek. Additionally, we designed a data screening pipeline that employs large language models to assess the logical consistency of outputs, ultimately yielding a high-quality distilled dataset.
@@ 2.2 Data Filtering
To ensure the quality and reliability of the distilled dataset, we implemented a rigorous data filtering pipeline consisting of multiple stages:
  • Model-Based Semantic Filtering: Employed LLM-based scoring (e.g., GPT-4) to assess:
    • Relevance: Whether answers properly address the questions.
    • Accuracy: Factual correctness based on medical knowledge.
    • Consistency: Logical coherence within the response.
  • Rule-Based Filtering:
    • Content consistency: Filtering incorrectly generated samples through character matching in JSON files or textbook captions.
    • Format screening: Filter out samples with overly short responses or excessive numbers of images, while also removing sensitive information
  • Expert Validation: Assess the rationality of dataset construction through medical expert evaluation, and eliminate unreasonable datasets.
Through the aforementioned data filtering process, we obtained a high-quality medical ultrasound dataset that covers textbooks and publicly available data, encompassing multiple anatomical regions and diverse diagnostic tasks.
@@
Dolphin v1.0 is built upon the Qwen2.5-VL model architecture, which represents an advanced vision-language foundation model developed by (Bai et al., 2025). The system enhances the original framework with a redesigned Vision Transformer (ViT) for optimized visual processing.
@@ 3.1 Training Recipe
The Dolphin v1.0 training pipeline follows a three-stage progressive framework, designed to integrate domain-specific knowledge, align with human preferences, and refine autonomous decision-making capabilities. The specific three stages are as follows: (1) Domain-Specialized Training; (2) Instruction-Driven Alignment; (3) Autonomous Reinforcement Refinement. Through two key phases - domain knowledge injection and instruction alignment - we developed the Dolphin V1.0 model. Building upon this foundation, we conducted reinforcement learning training using extensive ultrasound image-text data to obtain the Dolphin R1 model. Our subsequent research further explored the significance and impact of reinforcement learning on large-scale medical models.
@@ 3.1.1 Domain-Specialized Training
The goal of the Domain-Specialized Training is to inject ultrasound-specific knowledge into the large language model while preserving its inherent generalizability to the greatest extent possible. During this phase, the model is trained on extensive textbook-based ultrasound data and publicly available ultrasound datasets. The training encompasses ultrasound images of 15 major anatomical systems, with particular emphasis on developing fundamental capabilities in disease diagnosis, anatomical localization, and scan plane recognition. Notably, we transform lesion localization information into standardized directional descriptors (e.g., ”upper left quadrant”), thereby endowing the model with precise pathological assessment capabilities.
Recent studies conducted by (Chen et al., 2020), (Yang et al., 2025) have robustly evidenced that domain-specific fine-tuning of large language models with medical datasets may substantially compromise their core linguistic competencies, resulting in pronounced catastrophic forgetting phenomena. This empirical finding underscores the critical need for balanced adaptation strategies when specializing foundational models for medical applications.
During our model training process, we encountered the same issue. Therefore, when fine-tuning the model with ultrasound data using SFT (Supervised Fine-Tuning), we incorporated data from other domains, with particular emphasis on the model’s fundamental capabilities in the following aspects: reasoning, multi-turn dialogue, instruction-following , OCR detection, and mathematical reasoning. The introduction of general-domain data helped maintain these core competencies.
Interestingly, we observed that as the volume of both ultrasound-specific data and general-domain data increased proportionally, the model demonstrated simultaneous improvements in both ultrasound-related tasks and general-domain performance. This synergistic enhancement suggests that domain-specific specialization and general capability maintenance can be mutually reinforcing when balanced appropriately during training.
@@ 3.1.2 Instruction-Driven Alignment
At this stage, to prevent degradation of the model’s inherent capabilities, we adopted a relatively small learning rate and performed only 1 epoch of fine-tuning on the constructed Instruction-Driven Alignment dataset.
3.1.3 Autonomous Reinforcement Refinement.
Recent breakthroughs in reasoning models, such as OpenAI’s o-series ((Jaech et al., 2024)) and DeepSeek-R1 ((Guo et al., 2025)), have demonstrated remarkable performance in complex tasks, largely due to innovations in reinforcement learning (RL) techniques. Among these, Group Relative Policy Optimization (GRPO; (Shao et al., 2024)) has gained prominence for its efficiency in optimizing model behavior through verifiable reward signals. Unlike traditional supervised fine-tuning (SFT), which often suffers from answer overfitting and shortcut learning(Hao et al., 2025), GRPO encourages models to explore reasoning pathways autonomously, improving robustness and generalization—particularly critical in high-stakes domains like medicine(Pan et al., 2025).
For GRPO model training, verifiable results are of paramount importance. Each sample within our meticulously constructed UDP dataset is annotated with precise categories and measurement outcomes to guarantee data reliability and reproducibility. For each ultrasound task, we systematically generate corresponding questions based on the desired output format. The constructed dataset is categorized into two primary types:
  • General Responses: Open-ended answers requiring descriptive or analytical output.
  • Multiple-Choice Responses (A/B/C/D): Structured selections with rigorously controlled proportions to ensure balanced representation.
During the training process, we adopted the configuration proposed by (Guo et al., 2025), setting the reward function to consist of format reward and outcome reward, both assigned a value of 1.0. Through our exploration, we found that the model’s performance during the RL phase largely depends on the fundamental capabilities acquired in the SFT phase. The GRPO training approach primarily serves to activate the model’s inherent capabilities rather than inject new knowledge into it.
@@
To evaluate the ultrasound understanding capabilities of large vision-language models (LVLMs), we adopt U2-Bench (Le et al., 2025), a comprehensive benchmark specifical designed for ultrasound-related tasks. The experimental results presented in Table 1 provide a comprehensive evaluation of various large vision-language models (LVLMs) on the U2-Bench, a benchmark specifically designed to assess ultrasound-related multimodal understanding. Among the models evaluated, the Dolphin series—particularly Dolphin v1.0, Dolphin R1, and their smaller variant Dolphin v1.0-7B—demonstrate competitive performance across multiple tasks, with Dolphin R1 achieving the highest U2-Score of 0.5835 among all models tested.
@@ 4.1 Performance Overview
In comparison, Dolphin v1.0 shows slightly lower but still strong performance across these tasks, while Dolphin v1.0-7B, as a smaller variant, exhibits a noticeable performance drop, particularly in OD and KD, indicating the importance of model scale in capturing fine-grained visual features.
@@ 4.2 Impact of Model Scale
The comparison between Dolphin v1.0-7B and its larger 72B counterparts—Dolphin v1.0 and Dolphin R1—highlights the significant impact of model scale on multimodal understanding in ultrasound-related tasks. The 72B models consistently outperform the smaller 7B variant across classification, detection, and generation tasks, with the most notable improvements seen in Organ Detection (OD), Keypoint Detection (KD), and Caption Generation (CG), indicating that larger models are better suited to capture the visual complexity and semantic nuances present in medical ultrasound imaging.
Moreover, Dolphin R1 achieves a significantly improved U2-Score compared to Dolphin v1.0, demonstrating that advanced training strategies such as GRPO can effectively enhance model performance. This highlights the importance of both model scale and training methodology in pushing the boundaries of ultrasound image understanding.
@@ 4.3 Comparison Between Standard Mode and Deep Reasoning Mode
Figure 5: Comparison of Standard and Deep Reasoning models on DolphinV1.0 classification tasks, evaluated by accuracy, precision, recall, and F1-score.
As shown in Figure 5, we also evaluated the performance of our model, Dolphin R1, on disease diagnosis tasks under both standard mode and deep reasoning mode. The results demonstrate that, in the majority of tasks, enabling the deep reasoning mode leads to a notable improvement in diagnostic accuracy. Moreover, the deep reasoning mode offers a more thorough and detailed analytical process, providing interpretable and transparent reasoning trails. This enhanced interpretability not only supports more informed clinical decision-making but also aligns better with user preferences, as evidenced by the fact that physicians tend to favor the deep reasoning mode in practice.
@@ Conclusion
In this work, we introduced Dolphin v1.0 and Dolphin R1, advancing multimodal ultrasound understanding through large-scale data curation and progressive training strategies. The 200w-scale dataset provided the foundation for robust learning across heterogeneous tasks, while the integration of reinforcement learning in Dolphin R1 endowed the model with superior reasoning and interpretability. Our experimental results show that Dolphin R1 achieved a U2-score of 0.5835, surpassing all competitors by a large margin and doubling the performance of the second-best system, whereas Dolphin v1.0 established a strong baseline in standard mode. The comparative analysis highlights that deep reasoning not only improves quantitative accuracy but also enhances the interpretability of diagnostic processes, aligning more closely with physician preferences. Looking forward, this dual-track design of standard vs. deep reasoning modes provides a flexible paradigm for future medical AI systems, balancing efficiency, accuracy, and clinical usability.
@@
Contributors [^1] Alphabetical order.: Chi zhang, Chaoran Yan, Siya Liu, Xiaoyang Liu, Yalun Wu, Boyang Wang, Boyan Wang, Jiren Ren, Kaiwen Yan, Jinze Yu, Kaibing Hu, Henan Liu, Haoyun Zheng, Zhenyu Liu, Duo Zhang, Xiaoqing Guo

### v3

@@ 2.1.3 Ultrasound Domain Distillation Data
Following the methodology of Xu et al. (2023), we first constructed a set of ultrasound data question templates. The difficulty level of questions was enhanced through depth-wise augmentation, while their diversity was improved via breadth-wise expansion. All question categories were strictly standardized within the ultrasound medical domain, and we distilled ultrasound-specific data from both GPT and DeepSeek. Additionally, we designed a data screening pipeline that employs large language models to assess the logical consistency of outputs, ultimately yielding a high-quality distilled dataset.
@@
2.1.4 General Purpose Data
To enhance the model’s general capabilities while incorporating ultrasound-specific knowledge, we selected corresponding general-ability datasets based on different competency requirements. The included datasets were primarily utilized to enhance the following model capabilities: (1) Comprehension; (2) Conversational; (3) Deep reasoning; (4) Medical; (5) Image-text dialogue.
@@ 2.2 Data Filtering
To ensure the quality and reliability of the dataset, we designed a multi-stage filtering pipeline. In the first stage, model-based semantic filtering was applied, where LLM-based scoring (e.g., GPT-4) was used to evaluate the relevance of answers to the questions, their accuracy with respect to medical knowledge, and their logical consistency. This step helped remove outputs that were factually incorrect, irrelevant, or incoherent.
Next, we performed rule-based filtering. This included character matching to verify JSON integrity and caption consistency, as well as format screening to eliminate samples with overly short responses, excessive numbers of images, or the presence of sensitive information.
Finally, we conducted expert validation, in which board-certified physicians reviewed the remaining samples to assess the rationality of dataset construction and excluded any clinically unreasonable cases.
@@
Dolphin V1 is built upon the Qwen2.5-VL model architecture, which is an advanced large vision-language model developed by Bai et al. (2025). We enhanced the original framework with a redesigned Vision Transformer (ViT) for optimized visual processing.
@@ 3.1 Training Recipe
The Dolphin training pipeline follows a three-stage progressive framework, designed to integrate domain-specific knowledge, align with human preferences, and refine autonomous decision-making capabilities. The specific three stages are as follows: (1) Domain-Specialized Training; (2) Instruction-Driven Alignment; (3) Autonomous Reinforcement Refinement. Through two key phases - domain knowledge injection and instruction alignment - we developed the Dolphin V1 model. Building upon this foundation, we conducted reinforcement learning training using extensive ultrasound image-text data to obtain the Dolphin R1 model. Our subsequent research further explored the significance and impact of reinforcement learning on large-scale medical models.
@@ 3.1.1 Domain-Specialized Training
The goal of the Domain-Specialized Training is to inject ultrasound-specific knowledge into the large language model while preserving its inherent generalizability to the greatest extent possible. During this phase, the model is trained on extensive textbook-based ultrasound data and diagnostic-oriented ultrasound datasets. The training encompasses ultrasound images of 15 major anatomical systems, with particular emphasis on developing fundamental capabilities in disease diagnosis, anatomical localization, and standard plane recognition.
Recent studies (Chen et al., 2020; Yang et al., 2025) evidence the occurrence of catastrophic forgetting when fine-tuning LLMs on medical datasets, thereby compromising their core linguistic competencies, highlighting the need for balanced adaptation strategies in medical applications. Therefore, when fine-tuning the model with ultrasound data using SFT (Supervised Fine-Tuning), we incorporated data from other domains, with particular emphasis on the model’s fundamental capabilities in the following aspects: reasoning, multi-turn dialogue, instruction-following , OCR detection, and mathematical reasoning, which helped to maintain the core competencies of the original LLM. A figure illustrating the model’s performance on other benchmarks is included in Figure 6.
Figure 5: Model performance on General & Medical Bench
Interestingly, we observed that as the volume of both ultrasound-specific and general-domain data increased proportionally, the model demonstrated simultaneous improvements in both ultrasound-related tasks and general-domain performance. This synergistic enhancement suggests that domain-specific specialization and general capability maintenance can be mutually reinforcing when balanced appropriately during training.
@@ 3.1.2 Instruction-Driven Alignment
At this stage, to prevent degradation of the model’s inherent capabilities, we adopted a relatively small learning rate and performed only a few epoch of fine-tuning on the constructed Instruction-Driven Alignment dataset.
3.1.3 Autonomous Reinforcement Refinement
Recent advances in reasoning models, such as OpenAI’s o-series (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025), highlight the importance of reinforcement learning (RL) in enhancing diagnostic reasoning and interpretability. Unlike supervised fine-tuning (SFT), which is prone to overfitting and shortcut learning (Hao et al., 2025), RL enables models to autonomously explore reasoning pathways, thereby improving robustness and generalization in high-stakes domains such as medicine (Pan et al., 2025).
For ultrasound-specific reinforcement learning, we designed the Ultrasound Answer Reward (UAR), which evaluates model responses along two axes: format compliance and diagnostic accuracy. The reward is defined as:
$\text{UAR}(x,y)=0.5\times R_{\text{format}}(y)+0.5\times R_{\text{outcome}}(x,y),$ (1)
where $R_{\text{format}}(y)=1$ if the output follows the required structured format, and $R_{\text{outcome}}(x,y)=1$ if the response matches the ground-truth outcome. The outcome reward encompasses two types of verifiable results: (1) multiple-choice answers (A/B/C/D) for classification-style tasks, and (2) clinically accurate diagnostic conclusions for open-ended tasks, including disease identification, anatomical localization, and quantitative measurements.
This binary reward system ensures that responses are both structurally consistent and clinically valid. During reinforcement training, we observed that performance improvements strongly depended on the representational capabilities learned in the SFT phase; reinforcement learning primarily activated and refined these capabilities rather than introducing new medical knowledge.
@@ 3.1.3 Autonomous Reinforcement Refinement
To evaluate the ultrasound understanding capabilities of the trained model, we adopt U2-Bench (Le et al., 2025), a comprehensive benchmark specifical designed for ultrasound-related tasks. The experimental results presented in Table 1 provide a comprehensive evaluation of various MLLMs on the U2-Bench, a benchmark specifically designed to assess ultrasound-related multimodal understanding capability. Among the models evaluated, the Dolphin series, particularly Dolphin V1, Dolphin R1, and their smaller variant Dolphin V1-7B, demonstrate competitive performance across multiple tasks, with Dolphin R1 achieving the highest U2-Score of 0.5835 among all models tested.
@@
Figure 6: Comparison of Standard and Deep Reasoning models on DolphinV1 classification tasks, evaluated by accuracy, precision, recall, and F1-score.
@@ 4.1 Performance Overview
In comparison, Dolphin V1 shows slightly lower but still strong performance across these tasks, while Dolphin V1-7B, as a smaller variant, exhibits a noticeable performance drop, particularly in OD and KD, indicating the importance of model scale in capturing fine-grained visual features.
@@ 4.2 Impact of Model Scale
The comparison between Dolphin V1-7B and its larger 72B counterparts—Dolphin V1 and Dolphin R1—highlights the significant impact of model scale on multimodal understanding in ultrasound-related tasks. The 72B models consistently outperform the smaller 7B variant across classification, detection, and generation tasks, with the most notable improvements seen in Organ Detection (OD), Keypoint Detection (KD), and Caption Generation (CG), indicating that larger models are better suited to capture the visual complexity and semantic nuances present in medical ultrasound imaging.
Moreover, Dolphin R1 achieves a significantly improved U2-Score compared to Dolphin V1, demonstrating that advanced training strategies such as GRPO can effectively enhance model performance. This highlights the importance of both model scale and training methodology in pushing the boundaries of ultrasound image understanding.
@@ 4.3 Comparison Between Standard Mode and Deep Reasoning Mode
As shown in Figure 6, we also evaluated the performance of our model, Dolphin R1, on Disease Diagnosis (DD) tasks under both standard mode and deep reasoning mode. The results demonstrate that, in the majority of tasks, enabling the deep reasoning mode leads to a notable improvement in diagnostic accuracy. Moreover, the deep reasoning mode offers a more thorough and detailed analytical process, providing interpretable and transparent reasoning trails. This enhanced interpretability not only supports more informed clinical decision-making but also aligns better with user preferences, as evidenced by the fact that physicians tend to favor the deep reasoning mode in practice.
@@ Conclusion
In this work, we introduce Dolphin V1 and Dolphin R1, advancing multimodal ultrasound understanding through large-scale data curation and progressive training strategies. The 2-million-scale dataset provides the foundation for robust learning across heterogeneous tasks, while the integration of reinforcement learning in Dolphin R1 endows the model with superior reasoning and interpretability. Our experimental results show that Dolphin R1 achieved a U2-score of 0.5835, surpassing all competitors by a large margin. The comparative analysis highlights that deep reasoning not only improves quantitative accuracy but also enhances the interpretability of diagnostic processes, aligning more closely with physician preferences. Looking forward, this dual-track design of standard vs. deep reasoning modes provides a flexible paradigm for future medical AI systems.
@@
Contributors [^1] Alphabetical order.: Xiaoqing Guo, Kaibing Hu, Henan Liu, Siya Liu, Xiaoyang Liu, Zhenyu Liu, Jiren Ren, Boyan Wang, Boyang Wang, Yiyu Wang, Yalun Wu, Chaoran Yan, Kaiwen Yan, Jinze Yu, Chi Zhang, Duo Zhang, Haoyun Zheng
@@
Project Advisor: Jacques Souquet
Acknowledgement
We sincerely thank all the participating engineers and clinicians for their invaluable contributions throughout this project, including data labeling, validation, and quality checking. Their expertise and dedication ensured the clinical reliability of the datasets and the robustness of our evaluation.
We extend our special gratitude to Esonic Image for their generous support, including funding, provision of ultrasound devices for data collection, and coordination of medical experts to annotate datasets and provide continuous feedback during model development. Their collaboration was essential in bridging technical innovation with real-world clinical needs.